OpenStudy (sf_r):

Factor completely 3x^2+7x-6

OpenStudy (anonymous):

Try the quadratic formula

OpenStudy (sf_r):

Would this help I used the Asterisk method 

OpenStudy (calculusxy):

I don't really understand the Asterick method that you have posted. Do you know how to solve it using factor by grouping, quadratic formula, or completing the square? Although, I would recommend you to use factor by grouping to find the root of the quadratic.

OpenStudy (anonymous):

\[(3x - 2) \times ( x + 3) = 0\] \[3x -2 = 0 \] \[ x + 3 =0\] So answers should be x= 2/3 and x = -3

OpenStudy (sf_r):

The Quadratic formula would look like this, right